OpenStudy (anonymous):

Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received. x - 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive?

OpenStudy (anonymous):

x - 62,500 = 0.25y x-62,500=0.25*125,000

OpenStudy (anonymous):

Solve for x

OpenStudy (anonymous):

A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax, and x represents the original price of the item. C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale? Numerical Answers Expected!

OpenStudy (anonymous):

Did u solve for x?

OpenStudy (anonymous):

I dont know hwo to.. can you explain it?

OpenStudy (anonymous):

x - 62,500 = 0.25y x-62500=0.25*125000 x=31250+62500 x=93750

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

not reaslly haha I'm so dumb when it comes to math

OpenStudy (anonymous):

I guess u need to take extra algebra class